OpenStudy (anonymous):

Find an equation of the tangent line to the given curve at the given point. y= 2sinx at (pi/6 , 1)

OpenStudy (anonymous):

I found that y' = 2cosx y=mx+b b= y-mx = 1 - 2cosx (pi/6) am i right or wrong? what do i do next? i'm so off from the answer

OpenStudy (amistre64):

or in slope intercept form y = f'(a)x -f'(a)a+f(a)

OpenStudy (amistre64):

i see a typo in my first getup :) y - f(a) = f'(a) (x-a)

OpenStudy (anonymous):

where is b in here? or..is this some other formulas?

OpenStudy (anonymous):

cuz the only slope formula i know is y=mx+b..

OpenStudy (amistre64):

b = -f'(a)a+f(a)

OpenStudy (anonymous):

ohh.. i see..it's like the same so what i did was right..?

OpenStudy (amistre64):

b is a constant and can be derived from the point slope form of a line y - yo = m(x-xo) ; and solve for y y - yo = mx -mxo y = mx -mxo + yo

OpenStudy (amistre64):

you did good, lets check it tho

OpenStudy (anonymous):

but then my answer looks nothing similar to my textbook answer..

OpenStudy (amistre64):

y = 2sin(x) y' = 2cos(x) y = 2cos(30)x-2cos(30)pi/6+1

OpenStudy (amistre64):

cos 30 = sqrt(3)/2 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then y = sqrt(3)x -sqrt(3)pi/6 + 1

OpenStudy (amistre64):

how does the book look, cause theres different ways to simplify

OpenStudy (anonymous):

6 sqrt (3x) - 6y - pi sqrt (3) +6 =0

OpenStudy (amistre64):

ok; we can try to algebrate this to that

OpenStudy (amistre64):

y = sqrt(3)x -sqrt(3)pi/6 + 1 ; multiply it all by 6 6y = 6 sqrt(3)x -sqrt(3)pi + 6 ; bringit all to one side 0 = 6 sqrt(3)x -6y -sqrt(3)pi + 6

OpenStudy (anonymous):

wow..thanks! very helpful. it will take some time for my brain to process all this stuff :) but thanks! i get what u are saying :)

OpenStudy (amistre64):

good :)